OpenStudy (anonymous):

Select the best possible first step to solving the system by first eliminating the y variable. 4x - 5y = 4 3x + 2y = -20 Choose one answer. a. Multiply the first equation by -2 and multiply the second equation by 5. b. Multiply the first equation by 2 and multiply the second equation by -5. c. Multiply the first equation by 2 and multiply the second equation by 5. d. None of the above

OpenStudy (austinl):

C

OpenStudy (anonymous):

I really dont understand and any help is appriciated!

OpenStudy (austinl):

That will give you -10 and 10 which will cancel.

OpenStudy (anonymous):

Oh man, thank you so, so much!

OpenStudy (austinl):

You're welcome! If you want it explained more in depth I can if you like.

OpenStudy (anonymous):

Want to take a shot at another one and explain that one to me?

OpenStudy (austinl):

Sure.

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + y = 2 3x + 2y = 9

OpenStudy (austinl):

Ok, for this one we need to get x by itself in the first equation.

OpenStudy (austinl):

Which would be x=-y+2 right?

OpenStudy (anonymous):

Yes

OpenStudy (austinl):

so we have x=-y+2 and3x+2y=9 you put the x value into the second equation to get 3(-y+2)+2y=9

OpenStudy (anonymous):

okay so it's like figuring out a regular equation